# TV2TV: A Unified Framework for Interleaved Language and Video Generation

### v1

@@ 1 Introduction
We evaluate performance in two domains. First, we use video game data from Counter Strike; Global Offensive (CS:GO) curated and open-sourced by Pearce and Zhu (2022), which provides strongly correlated interleaved data via controller actions (represented as text) and resulting video gameplay. Controlled experiments on this data show that TV2TV substantially outperforms competing approaches in both visual quality – preferred $92\%$ of the time over a comparable T2V baseline – and controllability, with a 19 point improvement in fine-grained instruction-following accuracy compared to a “think-then-act” (Think2V) method which generates a detailed text action plan prior to generating video. Relatedly, action-conditioned world models (e.g., Bruce et al., 2024) can also generate game or synthetic video-world data. However, having them automatically generate videos without dense human control requires either a separate controller model (e.g., Raad et al., 2024) or a costly planning algorithm (e.g., Bar et al., 2025). Our TV2TV approach performs action generation and video-world generation flexibly and end-to-end.
@@
  • Controlled Experiments with Video Game Data (section 3) We validate this approach through controlled experiments with video game (CS:GO) data, demonstrating that TV2TV outperforms T2V and Think2V baselines in both visual quality and controllability. Specifically, in pairwise human evaluations, videos generated by TV2TV are preferred to those generated by a T2V baseline 92% of the time, and TV2TV shows a 19 point improvement in fine-grained instruction following accuracy compared to a Think2V baseline.
@@ 2 TV2TV
Figure 2: Overview of TV2TV architecture. TV2TV builds on the Transfusion (Zhou et al., 2024) modeling approach, jointly learning language modeling and video flow matching. TV2TV autoregressively generates interleaved chunks of video frames and text tokens, maintaining strict temporal causality: each token or chunk of frames can only attend to preceding tokens or chunks. TV2TV adopts a Mixture-of-Transformers (MoT) (Liang et al., 2024) architecture with dedicated towers for the video and text modalities.
@@
TV2TV adopts a Mixture-of-Transformers (MoT) (Liang et al., 2024) architecture with dedicated towers for the video and text modalities, enabling modality-specific processing while maintaining a global self-attention over the entire multimodal input sequence. Following Shi et al. (2024), we initialize the text tower with a pretrained language model. At inference, TV2TV flexibly alternates between text and video generation, factorizing the video generation into interleaved textual planning and video segment generation. See Figure 2 for an overview.
@@ 2.2 Architecture and optimization
where $\operatorname{mask}$ denotes a hybrid attention mask—applying a causal mask to the positions of text tokens and a block-causal mask to the positions of noisy and clean video tokens. An additional principle for masking is that noisy video tokens cannot be attended by any future tokens in the sequence. A global 1D RoPE is also applied here to all positions for all modalities.
@@ 4.2.2 Evaluation set-up
Figure 7: Evaluation of TV2TV trained on real world interleaved sports data vs. external video models We note that the external models we compare to were not specifically tuned for the sports domain and as such we do not expect them to perform well; however, we include them as out-of-domain baselines to obtain insights into TV2TV’s visual quality and prompt alignment relative to high-performing general-purpose models in this challenging domain. In blind pairwise human evaluations by external annotators, TV2TV outperforms all models on sports prompts in prompt alignment, real-world fidelity, and overall holistic preference. For visual quality, we find that TV2TV surpasses Cosmos2 variants, has similar performance as MAGI-1 variants, and underperforms compared to WAN 2.2 5B. 95% confidence intervals are shown.
@@
Figure 8: Evaluation of TV2TV trained on real world interleaved sports data vs. T2V and Think2V in a controlled setup. Compared to T2V, TV2TV shows stronger alignment and holistic preference, with similar visual quality and fidelity. When compared to Think2V, TV2TV shows a similar pattern of improved alignment and overall preference, though it is not statistically significant.
@@ 4.2.4 Qualitative analysis
Subsampled frames
t=0.0 t=1.19 t=2.38 t=3.62 t=4.81 t=6.06
Intervention 1 at t=1.56: “The man completes his golf swing, raising the club above his shoulder as he follows through. The golf ball is no longer visible, having been struck and sent flying forward. His body turns slightly to the left, maintaining balance after the powerful swing.”
Intervention 2 at t=1.56: “The man completes his golf swing, raising the club above his shoulder as he follows through. The camera pans to track the ball as it soars through the air.”
t=0.0 t=0.75 t=1.5 t=2.31 t=3.06 t=3.88
Intervention 1 at t=1.56: “The player in the white jersey takes control of the ball and runs towards the goal. He kicks the ball powerfully towards the net. The goalkeeper in the yellow jersey leaps to the left, extending his arms in an attempt to block the shot. The ball moves swiftly towards the goalpost as the goalkeeper’s jump reaches its peak.”
Intervention 2 at t=1.56: “The player in the red jersey intercepts the ball near the center of the field and starts dribbling towards the right side of the frame, moving away from the goal. He evades an approaching defender in a white jersey by maneuvering the ball skillfully. As he advances, other players adjust their positions, preparing for the next phase of play. The goalkeeper remains near the goal, observing the developing action.”
@@ 5 Related Work
Modeling long, sequential multimodal inputs, particularly video, reamains a critical challenge. Several approaches take an autoregressive approach to iteratively predict the next frame or block of frames conditioned on previous frames (Yin et al., 2024; Chen et al., 2025a; Agarwal et al., 2025; HaCohen et al., 2024; Lin et al., 2025; Cheng et al., 2025; Yuan et al., 2025), with notable work like MAGI-1 (Teng et al., 2025) focusing on high-fidelity AR video generation. However, these works, including MAGI-1, Cosmos (Agarwal et al., 2025), and Chen et al. (2025a), primarily focus on video generation and cannot generate interleaved text. Furthermore, Cosmos does not condition on the entire history, limiting long-term coherence. Meanwhile, token-based autoregressive generation—popularized by large language models (LLMs)—is utilized by models like VideoPoet (Kondratyuk et al., 2024), which treats video generation as a next-token prediction task and can straightforwardly leverage KV caching to improve generation efficiency.
To address the computational intensity and mitigate exposure bias during long-sequence generation, optimization techniques like diffusion forcing (Chen et al., 2025a) and self-forcing (Huang et al., 2025) are employed. Specifically, diffusion forcing introduced progressive noise to parallelize denoising across frames. To alleviate exposure bias in sequential training, a popular strategy is to inject a small amount of noise into clean data (Teng et al., 2025).
@@
Another line of work (Zheng et al., 2024; Peng et al., 2025b; Wan et al., 2025) follows the full-sequence paradigm where the video is treated as one large tensor and all frames are denoised simultaneously. However, this results in prohibitively large attention computations, making development of such methods prohibitive for long form generation.
@@ B.4 Human Evaluation Task
We include the evaluation questions answer by the annotators for the results discussed in section 4:

### v2

@@ 1 Introduction
We evaluate performance in two domains. First, we use video game data from Counter Strike; Global Offensive (CS:GO) curated and open-sourced by Pearce and Zhu (2022), which provides strongly correlated interleaved data via controller actions (represented as text) and resulting video gameplay. Controlled experiments on this data show that TV2TV substantially outperforms competing approaches in both visual quality – preferred $91\%$ of the time over a comparable T2V baseline – and controllability, with a 19 point improvement in fine-grained instruction-following accuracy compared to a “think-then-act” (Think2V) method which generates a detailed text action plan prior to generating video. Relatedly, action-conditioned world models (e.g., Bruce et al., 2024) can also generate game or synthetic video-world data. However, having them automatically generate videos without dense human control requires either a separate controller model (e.g., Raad et al., 2024) or a costly planning algorithm (e.g., Bar et al., 2025). Our TV2TV approach performs action generation and video-world generation flexibly and end-to-end.
@@
  • Controlled Experiments with Video Game Data (section 3) We validate this approach through controlled experiments with video game (CS:GO) data, demonstrating that TV2TV outperforms T2V and Think2V baselines in both visual quality and controllability. Specifically, in pairwise human evaluations, videos generated by TV2TV are preferred to those generated by a T2V baseline 91% of the time, and TV2TV shows a 19 point improvement in fine-grained instruction following accuracy compared to a Think2V baseline.
@@ 2 TV2TV
Figure 2: Overview of TV2TV architecture. TV2TV builds on the Transfusion (Zhou et al., 2024) modeling approach, jointly learning language modeling and video flow matching during training. TV2TV autoregressively generates interleaved chunks of video frames and text tokens, maintaining strict temporal causality: each token or chunk of frames can only attend to preceding tokens or chunks. TV2TV adopts a Mixture-of-Transformers (MoT) (Liang et al., 2024) architecture with dedicated towers for the video and text modalities.
@@
TV2TV adopts a Mixture-of-Transformers (MoT) (Liang et al., 2024) architecture with dedicated towers for the video and text modalities, enabling modality-specific processing while maintaining a global self-attention over the entire multimodal input sequence. Following Shi et al. (2024), we initialize the text tower with a pretrained language model. See Figure 2 for an overview of TV2TV’s architecture during training. At inference time, TV2TV flexibly alternates between text and video generation, factorizing the video generation into interleaved textual planning and video segment generation.
@@ 2.2 Architecture and optimization
where $\operatorname{mask}$ denotes a hybrid attention mask—applying a causal mask to the positions of text tokens and a block-causal mask to the positions of noisy and clean video tokens. An additional principle for masking is that noisy video tokens cannot be attended by any future tokens in the sequence. A global 1D RoPE is also applied here to all positions for all modalities, with duplicated position IDs for noisy and clean video tokens of the same frame chunk.
@@ 4.2.2 Evaluation set-up
Figure 7: Evaluation of TV2TV trained on interleaved sports data vs. external video models. We note that the external models we compare to were not specifically tuned for the sports domain and as such we do not expect them to perform well; however, we include them as out-of-domain baselines to obtain insights into TV2TV’s visual quality and prompt alignment relative to high-performing general-purpose models in this challenging domain. In blind pairwise human evaluations by external annotators, TV2TV outperforms all models on sports prompts in prompt alignment, real-world fidelity, and overall holistic preference. For visual quality, we find that TV2TV surpasses Cosmos2 variants, has similar performance as MAGI-1 variants, and underperforms compared to WAN 2.2 5B. 95% confidence intervals are shown.
@@
Figure 8: Evaluation of TV2TV trained on interleaved sports data vs. T2V and Think2V in a controlled setup. Compared to T2V, TV2TV shows stronger alignment and holistic preference, with similar visual quality and fidelity. When compared to Think2V, TV2TV shows a similar pattern of improved alignment and overall preference, though it is not statistically significant.
@@ 5 Related Work
Modeling long, sequential multimodal inputs, particularly video, remains a critical challenge. Several approaches take an autoregressive approach to iteratively predict the next frame or block of frames conditioned on previous frames (Yin et al., 2024; Chen et al., 2025a; Agarwal et al., 2025; HaCohen et al., 2024; Lin et al., 2025; Cheng et al., 2025; Yuan et al., 2025), with notable work like MAGI-1 (Teng et al., 2025) focusing on high-fidelity AR video generation. However, these works, including MAGI-1, Cosmos (Agarwal et al., 2025), and Chen et al. (2025a), primarily focus on video generation and cannot generate interleaved text. Furthermore, Cosmos does not condition on the entire history, limiting long-term coherence. Meanwhile, token-based autoregressive generation—popularized by large language models (LLMs)—is utilized by models like VideoPoet (Kondratyuk et al., 2024), which treats video generation as a next-token prediction task and can straightforwardly leverage KV caching to improve generation efficiency.
To address the computational intensity and mitigate exposure bias during long-sequence generation, optimization techniques like diffusion forcing (Chen et al., 2025a) and self-forcing (Huang et al., 2025) are employed. Specifically, diffusion forcing introduced progressive noise to parallelize denoising across frames. To alleviate exposure bias in sequential training, a popular strategy is to inject a small amount of noise into clean data (Teng et al., 2025). Our teacher-forcing setup that separately encodes noisy and clean data representations is also adopted in prior image generation work (Hu et al., 2024).
@@
Another line of work (Zheng et al., 2024; Peng et al., 2025b; Wan et al., 2025) follows the full-sequence paradigm where the video is treated as one large tensor and all frames are denoised simultaneously. However, this results in prohibitively large attention computations, making development of such methods prohibitive for long-form and interactive generation.
@@
Subsampled frames
t=0.0 t=1.19 t=2.38 t=3.62 t=4.81 t=6.06
Intervention 1 at t=1.56: The man completes his golf swing, raising the club above his shoulder as he follows through. The golf ball is no longer visible, having been struck and sent flying forward. His body turns slightly to the left, maintaining balance after the powerful swing.
Intervention 2 at t=1.56: The man completes his golf swing, raising the club above his shoulder as he follows through. The camera pans to track the ball as it soars through the air.
t=0.0 t=0.75 t=1.5 t=2.31 t=3.06 t=3.88
Intervention 1 at t=1.56: The player in the white jersey takes control of the ball and runs towards the goal. He kicks the ball powerfully towards the net. The goalkeeper in the yellow jersey leaps to the left, extending his arms in an attempt to block the shot. The ball moves swiftly towards the goalpost as the goalkeeper’s jump reaches its peak.
Intervention 2 at t=1.56: The player in the red jersey intercepts the ball near the center of the field and starts dribbling towards the right side of the frame, moving away from the goal. He evades an approaching defender in a white jersey by maneuvering the ball skillfully. As he advances, other players adjust their positions, preparing for the next phase of play. The goalkeeper remains near the goal, observing the developing action.
@@ B.4 Human Evaluation Task
We include the evaluation questions answered by the annotators for the results discussed in section 4: